OpenStudy (anonymous):

simplify 24- 10 and 5/12

OpenStudy (anonymous):

I don't really understand the format you put it in...

OpenStudy (anonymous):

You mean \(24 - 10\dfrac{5}{12}\)?

OpenStudy (anonymous):

yes i mean that

OpenStudy (e.cociuba):

its 14 and 5/12

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

yes rd42

OpenStudy (anonymous):

Let 24 = 23\(\dfrac{12}{12}\) 23 - 10 = 13 12/12 - 5/12 = 7/12 Is this clear?

OpenStudy (anonymous):

I mean e.c's answer is wrong.

OpenStudy (anonymous):

so 23 12/12 - 10 5/12 = 13 7/12

OpenStudy (anonymous):

Is this clear?

OpenStudy (anonymous):

yes it is clear thank you

hero (hero):

\(\large24\) \(\large-10 \frac{5}{12}\) \(\text{_______}\)

hero (hero):

\(\large23 \frac{12}{12}\) \(\large-10 \frac{5}{12}\) \(\text{_______}\) \(\large13 \frac{7}{12}\)

OpenStudy (anonymous):

Yeah, exactly what @Hero did.

hero (hero):

\[24 = 23 + 1 = 23 + \frac{12}{12} = 23 \frac{12}{12}\]